OpenStudy (anonymous):

Simplify the expression. 5C3

OpenStudy (anonymous):



Directrix (directrix):

C(5,3) means the number of ways to choose groups of 3 items from a set of 3. Order of selection does not matter. C(5,3) = 5! /[ (3! * (5 - 3) ! ] = 5! / ( 3! *2!) [ 5 * 4 * 3! ] /[ 3! * 2! ] = [ 5 * 4] / 2! = 20/2 = 10. --> Note that 2! = 2

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Any one else want to try this?

OpenStudy (alexwee123):